OpenStudy (anonymous):

WILL FAN AND MEDAL The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor (x - 7)2 represent? The square of the radius of the circular base of the tank The square of the height of the tank The area of the circular base of the tank The area of the curved sides of the tank

OpenStudy (cj49):

the square of the raidus of the circulare base of the tank

OpenStudy (phi):

radius of a cone-shaped tank is 4 feet less than its height. r is 4 less than h r = h-4 also they tell you height of the tank is (x - 3) . h= x-3 so r= h-4 becomes r= x-3 - 4 = x-7 so you know the radius is (x-7)

OpenStudy (phi):

also , if you see (for example) y^2 that means y*y , and people say "y squared" here you have (x-7)^2 and (x-7) is the radius. You can say it is "radius squared"